OpenStudy (anonymous):

How do I solve this? Find the solutions of g(x)=x^2+6x+1

geerky42 (geerky42):

Try to use quadratic formula.

OpenStudy (da_scienceman):

U can either use trail and error by guessing two numbers a and b such that a+b =1 and a*b06. The write down the factorization in the form (x+a)(x+b). Note that a,b or a and b can with be negative or both can be negative. Else you use the quadratic formula and substitute a=1, b=6, c=1 in to

OpenStudy (anonymous):

The formula is ax^2+bx+c, right?

OpenStudy (da_scienceman):

Its \[x _{1},x _{2}= \frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (da_scienceman):

I guess? that is the quadratice formula? and u write the resulting results as (x+x1)(x+x2). The plus and minus represent 2 roots.

OpenStudy (anonymous):

Not sure if I'm doing this right, but after plugging in the numbers to the formula I have 16 so far..

OpenStudy (da_scienceman):

Are u sure its g(x)=x^2+6x+1 and not either g(x)=x^2+6x-1 or x^2 -6x-1?

OpenStudy (anonymous):

Yes its g(x)=x^2+6x+1